# Continuous operators from spaces of Lipschitz functions

### v1

@@ 5.1. The Grothendieck property of spaces $\operatorname{Lip}_{0}(M)$
In this section we will provide several sufficient conditions for metric spaces $M$ implying that $\operatorname{Lip}_{0}(M)$ is not a Grothendieck space. Recall that a Banach space $X$ is Grothendieck (or has the Grothendieck property) if every weak* convergent sequence in the dual space $X^{*}$ is also weakly convergent. We refer the reader to [30] for an extensive survey on Grothendieck Banach spaces.
Typical examples of Grothendieck spaces are reflexive spaces, the space $\ell_{\infty}$, or, more generally, spaces $C(K)$ for $K$ extremally disconnected. On the other hand, spaces $C(K)$ for $K$ metric are never Grothendieck. Räbiger [55, Theorem 2.2] proved that a Banach space $X$ is a Grothendieck space if and only if $X^{*}$ is weakly sequentially complete and $X$ has no quotient isomorphic to $c_{0}$. Since for every compact space $K$ the dual space $C(K)^{*}$ is always weakly sequentially complete, Räbiger’s theorem, with an aid of the results of Cembranos [16] and Schachermayer [60], implies that a space $C(K)$ is not Grothendieck if and only if $C(K)$ has a quotient isomorphic to $c_{0}$ if and only if $C(K)$ contains a complemented copy of $c_{0}$. For spaces $\operatorname{Lip}_{0}(M)$ the situation is different—each space $\operatorname{Lip}_{0}(M)$ is isometrically isomorphic to the dual Banach space $\mathcal{F}(M)^{*}$, hence it cannot contain any complemented copy of $c_{0}$ (see e.g. [21, Theorem 2.4.15]), even though we know by Theorem 3.2 that it contains some copy of $c_{0}$. Consequently, to prove that a given space $\operatorname{Lip}_{0}(M)$ is not Grothendieck, using Räbiger’s criterion, we really need either to look for a quotient of $\operatorname{Lip}_{0}(M)$ isomorphic to $c_{0}$ or to show that the dual space $\operatorname{Lip}_{0}(M)^{*}$ is not weakly sequentially complete. To achieve the former, the results obtained in the previous section together with the standard fact that $\ell_{1}$ has a quotient isomorphic to $c_{0}$ might be useful.
@@
(1) Recall that a Banach space with property (V) is a Grothendieck space if and only if it does not contain a complemented copy of $c_{0}$, see e.g. [30, Proposition 3.1.13]. It follows that $E$ contains a complemented copy of $c_{0}$ and hence $E^{*}$ contains a complemented copy of $\ell_{1}$. This implies that $E^{*}$ is not Grothendieck.
(2) As $E$ is non-reflexive, $E^{*}$ is non-reflexive, too. By [36, Corollary 3.3.C], a non-reflexive space with property $(V^{*})$ contains a complemented copy of $\ell_{1}$ and hence is not Grothendieck.
@@
The space $\operatorname{Lip}_{0}(M)$ for infinite $M$ has never property $(V^{*})$, since, by [36, Proposition 3.3.E], property $(V^{*})$ is inherited by subspaces and, by Theorem 3.1, $\operatorname{Lip}_{0}(M)$ contains a copy of $\ell_{\infty}$, which does not have property $(V^{*})$ (since non-reflexive spaces with property $(V^{*})$ contain complemented copies of $\ell_{1}$, by [36, Corollary 3.3.C]).
@@
The results obtained in the first part of this section may be easily used to obtain further sufficient conditions for a Banach space $E$ implying that the space $\operatorname{Lip}_{0}(E)$ is not Grothendieck, thus completing Proposition 5.13.
@@
(1) follows from Proposition 5.2, (2) from Proposition 5.4, and (3) from Proposition 5.9. ∎
@@
By [55, Theorem 2.2], if the dual $E^{*}$ of a Banach space $E$ is not weakly sequentially complete, then $E$ is not a Grothendieck space. Using this theorem and the lifting property of separable Banach spaces, we make the following observation (cf. [13, Remark 3.6]).
@@
Since separable spaces have the lifting property (see [29, Theorem 3.1]), the Lipschitz-free space $\mathcal{F}(E)$ contains $E$ as a closed linear subspace (by [29, Theorem 2.12]), and hence $\operatorname{Lip}_{0}(E)^{*}\simeq\mathcal{F}(E)^{**}$ cannot be weakly sequentially complete. Consequently, $\operatorname{Lip}_{0}(E)$ does not have the Grothendieck property. ∎
Remark 5.23.
Recall that the James space $\mathcal{J}$ is an example of a non-reflexive quasi-reflexive separable infinite-dimensional Banach space (see [1, Section 3.4]). By the quasi-reflexivity, the dual space $\mathcal{J}^{*}$ is also separable, hence, by Proposition 5.13, $\operatorname{Lip}_{0}(\mathcal{J})$ is not Grothendieck. Note that also Proposition 5.22 immediately applies to $\mathcal{J}$. Indeed, again by the quasi-reflexivity, $\mathcal{J}$ does not contain any copies of $c_{0}$ or $\ell_{1}$. Consequently, by the non-reflexivity and Rosenthal’s $\ell_{1}$-theorem, $\mathcal{J}$ is not weakly sequentially complete.
@@
An infinite dimensional Banach space for which the space of Lipschitz functions does not have the Grothendieck property, but neither (1) nor (2) above is satisfied, is $\ell_{1}$. By [23, Proposition 79] the space $\ell_{1}$ is complemented in $\operatorname{Lip}_{0}(\ell_{1})$ and hence the latter space cannot have the Grothendieck property. Since $\ell_{1}$ has the Schur property, it is weakly sequentially complete, so (1) is not satisfied. In order to prove that (2) is not satisfied as well, we use that by the corollary to  [56, Theorem 5.1] there is no uniform and hence no bilipschitz embedding of $S_{c_{0}}$ into a stable Banach space and that a direct consequence of [56, Theorem 0.1] is that $\ell_{1}$ is stable [^1] The authors would like to thank Bunyamin Sari for pointing out Raynaud’s paper to them..
@@ 6. Open problems and final remarks
Theorem 4.8 provides several criteria implying that, for given infinite metric spaces $M$ and $N$, the spaces $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ are not homeomorphic. We are, however, not aware of any examples of spaces $M$ and $N$ such that $C_{p}(M)$ and $\mathcal{F}(M)_{w}$ are homeomorphic. Conditions (1)–(6) in Theorem 4.8 imply that if such spaces $M$ and $N$ do exist, then they both must be separable, $M$ must be non-scattered $\sigma$-compact and of cardinality $2^{\aleph_{0}}$, and $\mathcal{F}(N)$ cannot have the Schur property (which, by [3, Theorem C], is equivalent for the completion of $N$ to contain a bilipschitz copy of a subset of $\mathbb{R}$ of non-zero Lebesgue measure), nor it can be embeddable into a separable $L_{1}(\mu)$-space. Natural candidates could be then, e.g., $M=N=\mathbb{R}^{2}$.
@@
  • If $\operatorname{Lip}_{0}(\ell_{2})$ is a Grothendieck space, then we have an interesting situation, since by Remark 5.23 for the James space $\mathcal{J}$ the space $\operatorname{Lip}_{0}(\mathcal{J})$ is not Grothendieck and it is known that $\mathcal{J}$ is $\ell_{2}$-saturated, that is, each of its closed infinite-dimensional linear subspaces contains an isomorphic copy of the space $\ell_{2}$ (see [37, Lemma 1]).
@@
  • Recall that the space $\operatorname{Lip}_{0}(\mathbb{R}^{2})$ is isomorphic to the space $\operatorname{Lip}_{0}([0,1]^{2})$ (see [41, Corollary 3.5]), where the unit square $[0,1]^{2}$ carries the metric induced by the maximum norm on $\mathbb{R}^{2}$ and is an absolute Lipschitz retract. Consequently, if $\operatorname{Lip}_{0}(\mathbb{R}^{2})$ is not Grothendieck, then for every metric space $M$ containing a bilipschitz copy of $[0,1]^{2}$ the space $\operatorname{Lip}_{0}(M)$ does not have the Grothendieck property.
@@
For any $n\geq 1$, the space $\operatorname{Lip}_{0}(\mathbb{R}^{n})^{*}$ cannot have the Grothendieck property, since, by [20, Corollary 1.2], it contains $\mathcal{F}(\mathbb{R}^{n})$ as a complemented subspace and hence, by Corollary 3.3, also a complemented copy of $\ell_{1}$.
We finish the paper with the following important and well-known question (cf. [14, Introduction]). Note that a positive answer to Problem 6.3 or Problem 6.4 would provide a positive answer to Problem 6.6 as well.
@@
By the fact that for every Banach space $E$ the space $\operatorname{Lip}_{0}(E)$ contains a complemented copy of the dual space $E^{*}$, if for every two infinite-dimensional Banach spaces $E$ and $F$ of the same density the spaces $\operatorname{Lip}_{0}(E)$ and $\operatorname{Lip}_{0}(F)$ were isomorphic, then we would have the following interesting situation: given an infinite-dimensional Banach space $E$ of density $\kappa$, the space $\operatorname{Lip}_{0}(E)$ would contain a complemented copy of the dual space $F^{*}$ for every infinite-dimensional Banach space $F$ of density $\leq\kappa$. In particular, the space $\operatorname{Lip}_{0}(c_{0})$ would contain a complemented copy of the dual space $F^{*}$ for every separable Banach space $F$, which would provide a simple example of a Banach space with such a property and a separable predual. Let us note that such Banach spaces have been already constructed, though they are non-trivial, see e.g. [39, Theorem 1].
Acknowledgements. The first and the third named authors were supported by the Austrian Science Fund (FWF): I 4570-N. The third named author was also supported by the Austrian Science Fund (FWF): ESP 108-N.

### v3

@@ 5.1. The Grothendieck property of spaces $\operatorname{Lip}_{0}(M)$
Using results obtained in the first half of Section 5, we will now provide several sufficient conditions for metric spaces $M$ implying that $\operatorname{Lip}_{0}(M)$ is not a Grothendieck space. Recall that a Banach space $X$ is Grothendieck (or has the Grothendieck property) if every weak* convergent sequence in the dual space $X^{*}$ is also weakly convergent. We refer the reader to [32] for an extensive survey on Grothendieck Banach spaces.
Typical examples of Grothendieck spaces are reflexive spaces, the space $\ell_{\infty}$, or, more generally, spaces $C(K)$ for $K$ extremally disconnected. On the other hand, spaces $C(K)$ for $K$ metric are never Grothendieck. Räbiger [57, Theorem 2.2] proved that a Banach space $X$ is a Grothendieck space if and only if $X^{*}$ is weakly sequentially complete and $X$ has no quotient isomorphic to $c_{0}$. Since for every compact space $K$ the dual space $C(K)^{*}$ is always weakly sequentially complete, Räbiger’s theorem, with an aid of the results of Cembranos [17] and Schachermayer [62], implies that a space $C(K)$ is not Grothendieck if and only if $C(K)$ has a quotient isomorphic to $c_{0}$ if and only if $C(K)$ contains a complemented copy of $c_{0}$. For spaces $\operatorname{Lip}_{0}(M)$ the situation is different—each space $\operatorname{Lip}_{0}(M)$ is isometrically isomorphic to the dual Banach space $\mathcal{F}(M)^{*}$, hence it cannot contain any complemented copy of $c_{0}$ (see e.g. [23, Theorem 2.4.15]), even though we know by Theorem 3.2 that it contains some copy of $c_{0}$. Consequently, to prove that a given space $\operatorname{Lip}_{0}(M)$ is not Grothendieck, using Räbiger’s criterion, we really need either to look for a quotient of $\operatorname{Lip}_{0}(M)$ isomorphic to $c_{0}$ or to show that the dual space $\operatorname{Lip}_{0}(M)^{*}$ is not weakly sequentially complete. To achieve the former, the results obtained in the previous section together with the standard fact that $\ell_{1}$ has a quotient isomorphic to $c_{0}$ might be useful.
@@
(1) Recall that a Banach space with property (V) is a Grothendieck space if and only if it does not contain a complemented copy of $c_{0}$, see e.g. [32, Proposition 3.1.13]. It follows that $E$ contains a complemented copy of $c_{0}$ and hence $E^{*}$ contains a complemented copy of $\ell_{1}$. This implies that $E^{*}$ is not Grothendieck.
(2) As $E$ is non-reflexive, $E^{*}$ is non-reflexive, too. By [38, Corollary 3.3.C], a non-reflexive space with property $(V^{*})$ contains a complemented copy of $\ell_{1}$ and hence is not Grothendieck.
@@
The space $\operatorname{Lip}_{0}(M)$ for infinite $M$ has never property $(V^{*})$, since, by [38, Proposition 3.3.E], property $(V^{*})$ is inherited by subspaces and, by Theorem 3.1, $\operatorname{Lip}_{0}(M)$ contains a copy of $\ell_{\infty}$, which does not have property $(V^{*})$ (since non-reflexive spaces with property $(V^{*})$ contain complemented copies of $\ell_{1}$, by [38, Corollary 3.3.C]).
@@
Results obtained in the first part of this section may be quickly used to obtain further sufficient conditions for a Banach space $E$ implying that the space $\operatorname{Lip}_{0}(E)$ is not Grothendieck, thus completing Proposition 5.14.
@@
(1) follows from Proposition 5.2, (2) from Lemma 5.4, and (3) from Corollary 5.9. ∎
@@
By [57, Theorem 2.2], if the dual $E^{*}$ of a Banach space $E$ is not weakly sequentially complete, then $E$ is not a Grothendieck space. Using this theorem and the lifting property of separable Banach spaces, we make the following observation (cf. [13, Remark 3.6]), which can be applied to the James space $\mathcal{J}$.
@@
Since separable spaces have the lifting property (see [31, Theorem 3.1]), the Lipschitz-free space $\mathcal{F}(E)$ contains $E$ as a closed linear subspace (by [31, Theorem 2.12]), and hence $\operatorname{Lip}_{0}(E)^{*}\simeq\mathcal{F}(E)^{**}$ cannot be weakly sequentially complete. Consequently, $\operatorname{Lip}_{0}(E)$ does not have the Grothendieck property. ∎
Recall that the James space $\mathcal{J}$ is an example of a non-reflexive quasi-reflexive separable infinite-dimensional Banach space (see [1, Section 3.4]).
Corollary 5.24.
$\operatorname{Lip}_{0}(\mathcal{J})$ is not Grothendieck.
Proof.
The statement can be justified twofold. First, by the quasi-reflexivity, the dual space $\mathcal{J}^{*}$ is also separable, hence, by Proposition 5.14, $\operatorname{Lip}_{0}(\mathcal{J})$ is not Grothendieck. Second, again by the quasi-reflexivity, $\mathcal{J}$ does not contain any copies of $c_{0}$ or $\ell_{1}$. Consequently, by the non-reflexivity and Rosenthal’s $\ell_{1}$-theorem, $\mathcal{J}$ is not weakly sequentially complete. Thus, Proposition 5.23 applies to $\mathcal{J}$, too. ∎
@@
An infinite dimensional Banach space for which the space of Lipschitz functions does not have the Grothendieck property, but neither (1) nor (2) above is satisfied, is $\ell_{1}$. By [24, Proposition 79] the space $\ell_{1}$ is complemented in $\operatorname{Lip}_{0}(\ell_{1})$ and hence the latter space cannot have the Grothendieck property. Since $\ell_{1}$ has the Schur property, it is weakly sequentially complete, so (1) is not satisfied. In order to prove that (2) is not satisfied as well, we use that by the corollary to  [58, Theorem 5.1] there is no uniform and hence no bilipschitz embedding of $S_{c_{0}}$ into a stable Banach space and that a direct consequence of [58, Theorem 0.1] is that $\ell_{1}$ is stable [^1] The authors would like to thank Bunyamin Sari for pointing out Raynaud’s paper to them..
@@ 6. Open problems and final remarks
Theorem 4.8 provides several criteria implying that, for given infinite metric spaces $M$ and $N$, the spaces $C_{p}(M)$ and $\mathcal{F}(N)_{w}$ are not homeomorphic. We are, however, not aware of any examples of spaces $M$ and $N$ such that $C_{p}(M)$ and $\mathcal{F}(M)_{w}$ are homeomorphic. Conditions (1)–(6) in Theorem 4.8 imply that if such spaces $M$ and $N$ do exist, then they both must be separable, $M$ must be non-scattered $\sigma$-compact and of cardinality $2^{\aleph_{0}}$, and $\mathcal{F}(N)$ cannot have the Schur property (which, by [2, Theorem C], is equivalent for the completion of $N$ to contain a bilipschitz copy of a subset of $\mathbb{R}$ of non-zero Lebesgue measure), nor it can be embeddable into a separable $L_{1}(\mu)$-space. Natural candidates could be then, e.g., $M=N=\mathbb{R}^{2}$.
@@
  • If $\operatorname{Lip}_{0}(\ell_{2})$ is a Grothendieck space, then we have an interesting situation, since by Corollary 5.24 for the James space $\mathcal{J}$ the space $\operatorname{Lip}_{0}(\mathcal{J})$ is not Grothendieck and it is known that $\mathcal{J}$ is $\ell_{2}$-saturated, that is, each of its closed infinite-dimensional linear subspaces contains an isomorphic copy of the space $\ell_{2}$ (see [39, Lemma 1]).
@@
  • Recall that the space $\operatorname{Lip}_{0}(\mathbb{R}^{2})$ is isomorphic to the space $\operatorname{Lip}_{0}([0,1]^{2})$ (see [44, Corollary 3.5]), where the unit square $[0,1]^{2}$ carries the metric induced by the maximum norm on $\mathbb{R}^{2}$ and is an absolute Lipschitz retract. Consequently, if $\operatorname{Lip}_{0}(\mathbb{R}^{2})$ is not Grothendieck, then for every metric space $M$ containing a bilipschitz copy of $[0,1]^{2}$ the space $\operatorname{Lip}_{0}(M)$ does not have the Grothendieck property.
@@
For any $n\geq 1$, the space $\operatorname{Lip}_{0}(\mathbb{R}^{n})^{*}$ cannot have the Grothendieck property, since, by [22, Corollary 1.2], it contains $\mathcal{F}(\mathbb{R}^{n})$ as a complemented subspace and hence, by Corollary 3.3, also a complemented copy of $\ell_{1}$.
We finish the paper with the following important and well-known question (cf. [15, Introduction]). Note that a positive answer to Problem 6.3 or Problem 6.4 would provide a positive answer to Problem 6.6 as well.
@@
By the fact that for every Banach space $E$ the space $\operatorname{Lip}_{0}(E)$ contains a complemented copy of the dual space $E^{*}$, if for every two infinite-dimensional Banach spaces $E$ and $F$ of the same density the spaces $\operatorname{Lip}_{0}(E)$ and $\operatorname{Lip}_{0}(F)$ were isomorphic, then we would have the following interesting situation: given an infinite-dimensional Banach space $E$ of density $\kappa$, the space $\operatorname{Lip}_{0}(E)$ would contain a complemented copy of the dual space $F^{*}$ for every infinite-dimensional Banach space $F$ of density $\leq\kappa$. In particular, the space $\operatorname{Lip}_{0}(c_{0})$ would contain a complemented copy of the dual space $F^{*}$ for every separable Banach space $F$, which would provide a simple example of a Banach space with such a property and a separable predual. Let us note that such Banach spaces have been already constructed, though they are non-trivial, see e.g. [41, Theorem 1].
Acknowledgements
The authors wish to thank the anonymous referee for carefully reading the paper and for a number of valuable suggestions.
The first and the third named authors were supported by the Austrian Science Fund (FWF): I 4570-N. The third named author was also supported by the Austrian Science Fund (FWF): ESP 108-N.
This research was funded in whole or in part by the Austrian Science Fund (FWF), grant DOI: 10.55776/ESP108, 10.55776/I4570. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.